OpenStudy (inowalst):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (sierra!):

What do you think?

OpenStudy (inowalst):

I have no idea, that's why I asked.

OpenStudy (sierra!):

OK lets start out simple

OpenStudy (sierra!):

HR = Horse and HM= Human

OpenStudy (sierra!):

ok?

OpenStudy (inowalst):

Okay.

OpenStudy (sierra!):

So HM + HR =?

OpenStudy (inowalst):

74

OpenStudy (sierra!):

Yup!

OpenStudy (sierra!):

So 2HM + 4HR =?

OpenStudy (inowalst):

196?

OpenStudy (sierra!):

Correct!

OpenStudy (sierra!):

So (2HM + 4HR) - (2 HM + 2HR) =?

OpenStudy (inowalst):

Okay, give me a minute.

OpenStudy (sierra!):

Yup! take all the time you need :)

OpenStudy (inowalst):

I'm not so sure here. :/

OpenStudy (sierra!):

What its saying is 196 - 148=?

OpenStudy (inowalst):

48

OpenStudy (sierra!):

Yup!

OpenStudy (sierra!):

So 2HR = 48 and HR = 24

OpenStudy (here_to_help15):

h + p=74 4h + 2p=196 (horses have 4 legs people have 2) p=74 - h 4h + 2(74 - h)=196 (plug (74 - h) in for p, then solve) 4h + 148 - 2h=196 2h=48 h=24 (which means p=50) There are 24 horses and 50 humans!

OpenStudy (sierra!):

So HM + (24) =?

OpenStudy (sierra!):

No direct answers i understand you showed the work but he/she has to get the answers theirself

OpenStudy (sierra!):

please read the code of conduct

OpenStudy (here_to_help15):

I have

OpenStudy (sierra!):

Then you should know what to do instead of giving the answer

OpenStudy (inowalst):

@sierra! 74?

OpenStudy (sierra!):

Yes!

OpenStudy (sierra!):

And lastly HM = 74 - 24=?

OpenStudy (inowalst):

50

OpenStudy (sierra!):

Yes!

OpenStudy (here_to_help15):